Part A Find the amount of heat that must be extracted from 1.0kg of steam at...

Part A

Find the amount of heat that must be extracted from 1.0kg of steam at 150?C to convert it to ice at 0.0 ?C.

Part B

What speed would this 1.0-kg block of ice have if its translational kinetic energy were equal to the thermal energy calculated in part A?
Answer #1

A)

mass , m = 1 Kg

heat removed = m*(Sice * 50 + Lv + Swater* 100 + Lf)

heat removed = 1 *(2108 *50 + 2270 *10^3 + 4186 * 100 + 334*10^3 )

heat removed = 3128000

heat removed is 3128000 J

B)

Using formula for kinetic energy

0.5 * 1 * u^2 = 3128000

u = 2501.2 m/s

the speed of block is 2501.2 m/s